### v1

# Mind the Confidence Gap: Overconfidence, Calibration, and Distractor Effects in Large Language Models

Abstract
Large Language Models (LLMs) demonstrate impressive performance across diverse tasks, yet confidence calibration remains a challenge. Miscalibration—where models are overconfident or underconfident—poses risks, particularly in high-stakes applications. This paper presents an empirical study on LLM calibration, examining how model size, distractors, and question types affect confidence alignment. We introduce an evaluation framework to measure overconfidence and investigate whether multiple-choice formats mitigate or worsen miscalibration. Our findings show that while larger models (e.g., GPT-4o) are better calibrated overall, they are more prone to distraction, whereas smaller models benefit more from answer choices but struggle with uncertainty estimation. Unlike prior work, which primarily reports miscalibration trends, we provide actionable insights into failure modes and conditions that worsen overconfidence. These findings highlight the need for calibration-aware interventions and improved uncertainty estimation methods. To ensure reproducibility, we provide our code and dataset [^1] https://github.com/prateekchhikara/llms-calibration.
Mind the Confidence Gap: Overconfidence, Calibration, and Distractor Effects in Large Language Models
Prateek Chhikara University of Southern California, USA pchhikar@usc.edu
1 Introduction
Large Language Models (LLMs) have transformed natural language understanding, achieving state-of-the-art performance across diverse applications, from conversational AI Skjuve et al. (2024); Zhang (2024) to scientific discovery Kumar (2024). Their versatility has also facilitated advancements in multimodal learning, where language models integrate with vision-based systems to enhance tasks such as automated content generation and decision support Zhang et al. (2023); Chhikara et al. (2024). However, as LLMs are increasingly deployed in high-stakes domains such as healthcare, finance, and law, critical concerns arise regarding calibration—the alignment between model confidence and actual correctness Dhuliawala et al. (2023). When LLMs express unwarranted confidence in incorrect predictions, they risk misleading users, spreading misinformation, and reducing trust and reliability in AI-driven systems Chen et al. (2023); Zhang et al. (2024).
Figure 1: LLM generating an incorrect answer with high confidence.
Figure 1 illustrates this issue: when asked “Who received the IEEE Frank Rosenblatt Award in 2010?”, an LLM incorrectly responds with “Geoffrey Hinton”, assigning a 93% confidence score, despite correct answer being “Michio Sugeno”. This overconfidence is especially problematic because users tend to equate high-confidence outputs with reliability Liu et al. (2024). The right-hand side of Figure 1 further demonstrates the broader miscalibration problem: while a well-calibrated model should align confidence with accuracy, real-world LLMs frequently overestimate their correctness.
Although calibration errors in neural networks are well-documented Guo et al. (2017), the calibration behavior of LLMs remains poorly understood, particularly in response to fine-tuning techniques such as Reinforcement Learning from Human Feedback (RLHF) Leng et al. (2024); Li et al. (2024). Prior work Zhu et al. (2023); Geng et al. (2024) has noted that RLHF can amplify overconfidence, but the extent and specific conditions under which this occurs remain largely unexplored. Key open questions include: (i) Do bigger models exhibit better calibration, or does scale introduce new sources of miscalibration? and (ii) Can structured input formats, such as multiple-choice answer choices, improve confidence alignment, or do they introduce additional biases? Addressing these gaps is crucial for ensuring the reliable deployment of LLMs in decision-critical applications.
This paper presents an empirical study of LLM calibration, focusing on three key objectives: (i) Quantifying Overconfidence: We systematically measure miscalibration across multiple LLMs and analyze the extent of overconfidence. (ii) Evaluating the Role of Answer Choices: We assess whether providing structured answer options (e.g., multiple-choice format) improves or worsens accuracy and calibration. (iii) Identifying Failure Modes: We categorize question types where LLMs exhibit the highest miscalibration, revealing systematic weaknesses in uncertainty estimation.
Unlike prior work Ulmer et al. (2024); Kapoor et al. (2024); Jiang et al. (2021) that broadly characterizes miscalibration, our study pinpoints specific conditions that worsen overconfidence, offering an actionable framework for improving LLM reliability. Our findings emphasize the need for task-specific confidence adjustments and structured uncertainty modeling to enhance trust in real-world deployments. Our study reveals that miscalibration is not a uniform problem; its severity varies based on model scale and input structure, with larger models gaining significant calibration improvements when provided with distractors, whereas smaller models struggle with meaningful uncertainty estimation.
2 Experimental Setup
Evaluation Dataset:
We use the SimpleQA dataset Wei et al. (2024), which provides a reliable benchmark for evaluating LLM factual accuracy and calibration. Comprising short, fact-seeking queries with clearly defined correct answers, SimpleQA enables precise measurement of model confidence and alignment with factual correctness. Its high-quality annotations, verified by multiple independent AI trainers, ensure accuracy and unambiguity, making it well-suited for calibration assessment. The dataset contains 4326 question-answer pairs. For each pair, we used GPT-4o-mini to generate three distractors—factually incorrect yet contextually plausible answers that matched the answer type (e.g., dates for date-based questions), remained distinct, and maintained similar specificity. Additional details on the prompt are provided in Appendix A.
Selected LLMs:
We select models from OpenAI [^2] https://openai.com and GroqCloud [^3] https://groq.com. From OpenAI, we include GPT-4o, GPT-4o-mini, and GPT-4-turbo Hurst et al. (2024), all large-scale models with billions of parameters. From GroqCloud, we select smaller, efficiency-optimized models: LLaMA-3-8B-8192 Dubey et al. (2024), LLaMA3.1-8B-instant, and Gemma2-9B-it Team et al. (2024). More details about the selected models are in Appendix C.
Table 1: Performance metrics of LLMs in the Normal ($\mathcal{N}$) and Distractor ($\mathcal{D}$) settings, including accuracy (correct), NOT_ATTEMPTED (na), ECE, and the number of helped ($\mathcal{D}_{helped}$) and harmed ($\mathcal{D}_{harmed}$) instances.
LLMs $\mathcal{N}_{correct}$ $\mathcal{N}_{na}$ $\mathcal{N}_{ECE}$ $\mathcal{D}_{correct}$ $\mathcal{D}_{na}$ $\mathcal{D}_{ECE}$ $\mathcal{D}_{helped}$ $\mathcal{D}_{harmed}$
GPT-4o-mini [Uncaptioned image] 8.46% 6.80% 0.750 47.43% 0.02% 0.320 1644 (93.78%) 109 (6.22%)
GPT-4-turbo [Uncaptioned image] 20.37% 6.17% 0.612 65.40% 0% 0.165 1877 (95.86%) 81 (4.14%)
GPT-4o [Uncaptioned image] 35.14% 7.88% 0.450 73.42% 0.02% 0.037 1569 (91.97%) 137 (8.03%)
LLaMA3.1-8b-instant [Uncaptioned image] 5.58% 18.78% 0.799 44.64% 0.12% 0.367 1355 (95.29%) 67 (4.71%)
LLaMA 3-8B-8192 [Uncaptioned image] 4.79% 21.20% 0.810 44.01% 0.55% 0.361 1382 (95.57%) 62 (4.43%)
Gemma2-9B-it [Uncaptioned image] 5.48% 33.29% 0.799 45.58% 1.78% 0.367 1143 (94.38%) 68 (5.62%)
Figure 2: Reliability diagrams (RDs) showing calibration performance in $\mathcal{N}$ ($\bullet$) and $\mathcal{D}$ ($\bullet$) settings. The numbers on top of bars represent the number of correctly predicted instances. (y-axis: actual accuracy, x-axis: predicted confidence)
Evaluation Criteria
Following prior work, we use GPT-4o-mini as an LLM-based judge to classify responses as CORRECT, INCORRECT, or NOT_ATTEMPTED Packer et al. (2023); Wei et al. (2024). A response is CORRECT if it fully captures the gold target’s key information without contradiction, allowing minor variations in wording, order, or hedging. It is INCORRECT if it contains factual errors, contradictions, or misleading speculation, even if hedged. NOT_ATTEMPTED applies when a response lacks essential information without introducing errors, including vague or evasive answers. We experiment with using the same LLM for both prediction and judgment, finding that smaller LLM judges often misclassify responses or hesitate to assign NOT_ATTEMPTED when no valid answer is generated. Manual inspection confirms these issues, and further details are provided in the Appendix B.
Evaluation Methods:
The standard LLM approach ($\mathcal{N}$) generates answers based only on the given question and a predefined prompt, without external knowledge. We introduce an alternative distractor method ($\mathcal{D}$), where the model receives the question along with the correct answer and three incorrect options (shuffled). This setup enables us to evaluate whether additional context enhances calibration, reinforces overconfidence, or simply encourages random guessing from the provided options. Details on prompts are in the Appendix A.
Evaluation Metrics:
To evaluate performance, we measure correctly answered questions for both variations ($\mathcal{N}$ and $\mathcal{D}$). For calibration assessment, we use Expected Calibration Error (ECE) to quantify the misalignment between a model’s predicted confidence and actual accuracy. A well-calibrated model produces confidence estimates that closely match its true correctness, with an ECE of zero indicating perfect calibration. Following Naeini et al. (2015), we compute ECE using empirical binning (bin size 0.1) to ensure a robust measurement of miscalibration. Additionally, we introduce two metrics, $\mathcal{D}_{helped}$ and $\mathcal{D}_{harmed}$, which measure the number of instances where adding options either improved or worsened the model’s performance.
3 Results Analysis and Discussion
Performance across LLMs
Results indicate a clear performance gap between bigger and smaller LLMs, with bigger models (e.g., GPT-4o) significantly outperforming smaller ones in open-ended question answering (Table 1). However, when provided with structured answer choices, smaller models exhibit disproportionately higher gains, while bigger models experience relative increase in distractor-induced errors ($\mathcal{D}_{harmed}$). This suggests that bigger models rely more heavily on associative recall and self-generated confidence heuristics, which can be misleading when plausible but incorrect distractors are introduced.
For instance, GPT-4o’s accuracy jumps from 35.14% ($\mathcal{N}$) to 73.42% ($\mathcal{D}$), while LLaMA3-8B exhibits an even more dramatic improvement from 4.79% to 44.01%. The fact that smaller models show a greater relative benefit from answer choices suggests that their decision-making relies more on structured elimination strategies rather than independent fact retrieval. This finding highlights a key consideration for practical deployment: while bigger models can be more self-reliant, their confidence in unaided settings does not necessarily translate into robustness when faced with misleading choices.
Figure 3: Performance (correct) of LLMs across different question types in both $\mathcal{N}$ ($\bullet$) and $\mathcal{D}$ ($\bullet$) settings.
Calibration and Confidence Patterns
A well-calibrated model should align its confidence scores with actual accuracy—meaning that if a model expresses 80% confidence in an answer, it should be correct roughly 80% of the time. Our study reveals that bigger models, while better calibrated overall, still exhibit systematic overconfidence in open-ended tasks. The ECE for GPT-4o decreases significantly in the $\mathcal{D}$ setting, demonstrating that answer choices improve confidence alignment by providing external constraints. However, smaller models, while benefiting from options, still fail to estimate their uncertainty meaningfully, as evidenced by consistently high ECE values.
Figure 2 shows that calibration behavior varies significantly across models. While GPT-4o is relatively well-calibrated at higher confidence levels (70–100%), minor overconfidence remains. In contrast, smaller models frequently overestimate their accuracy, assigning high confidence to incorrect answers, particularly in the 80–100% range. Additionally, all LLMs at lower confidence levels (20–40%) show inconsistencies in uncertainty estimation, sometimes underestimating correctness but not necessarily engaging in random guessing. This distinction has significant implications in real-world applications: in AI-driven tutoring or medical diagnosis, overconfidence in incorrect answers can lead to misleading recommendations. Calibration-aware interventions—such as confidence regularization or explicit uncertainty weighting—are necessary to mitigate these risks.
Performance Across Question Types
To analyze calibration weaknesses, we evaluate performance across different question types (Figure 3). Person-based queries are most challenging, likely due to name ambiguities and inherent variability in names, overlapping roles, and contextual dependencies that require deeper reasoning beyond surface-level pattern matching. LLMs frequently confuse historical figures with similar names, but providing structured answer choices significantly improves accuracy, suggesting that explicit disambiguation helps mitigate uncertainty in this category. In contrast, place-based queries exhibit relatively strong performance across both settings, indicating that geographic knowledge is well-represented in pretraining. However, calibration improvements vary (Table 2): the person category sees highest relative ECE drop (69%), while place category shows the lowest (50%). This suggests that structured choices help correct overconfidence in ambiguous queries but offer limited calibration gains when models already retrieve knowledge with high confidence. These results demonstrate that miscalibration depends on both task framing and knowledge representation, not just model scale. While structured reasoning improves confidence alignment in person-based queries, factual retrieval tasks like place-based questions may require alternative calibration strategies to prevent persistent overconfidence.
Table 2: ECE comparison across question types.
Date Number Person Place
$\mathcal{N}$ $\mathcal{D}$ $\mathcal{N}$ $\mathcal{D}$ $\mathcal{N}$ $\mathcal{D}$ $\mathcal{N}$ $\mathcal{D}$
4o-mini 0.77 0.32 0.73 0.35 0.76 0.25 0.73 0.42
4-turbo 0.62 0.23 0.65 0.26 0.62 0.03 0.52 0.19
4o 0.41 0.05 0.53 0.13 0.51 0.07 0.35 0.04
LLaMA3.1 0.84 0.37 0.82 0.38 0.80 0.31 0.73 0.40
LLaMA3 0.85 0.35 0.82 0.36 0.81 0.33 0.73 0.45
Gemma2 0.80 0.33 0.78 0.42 0.84 0.33 0.76 0.41
mean 0.72 0.28 0.72 0.32 0.72 0.22 0.64 0.32
Implications for Model Deployment
Our findings indicate that calibration errors are not uniform and vary based on both model size and input structure. The susceptibility of bigger models to distractors suggests that prompt engineering strategies should include explicit calibration constraints—for example, by encouraging models to justify their confidence levels rather than relying on raw probability scores. For smaller models, answer ranking or re-scoring techniques may provide more robust confidence adjustments, ensuring that they do not simply default to high-confidence selections when uncertain. Future research should explore task-specific calibration techniques rather than applying one-size-fits-all solutions, particularly for applications where overconfidence poses significant risks (e.g., legal and medical AI systems).
4 Conclusion
This study provides an investigation of calibration in LLMs, revealing key trade-offs between model size, answer format, and confidence alignment. While bigger models exhibit better calibration, they remain susceptible to distraction, whereas smaller models show dramatic performance improvements with structured answer choices but fail to meaningfully estimate uncertainty. Unlike prior work that broadly characterizes LLM miscalibration, our study pinpoints conditions under which calibration errors arise, offering a framework for mitigating overconfidence in real-world applications. Future work should explore calibration-aware fine-tuning strategies and uncertainty-aware response generation to enhance LLM reliability, particularly in high-stakes domains.
5 Limitations
Our study is limited by its focus on a small set of LLMs, excluding domain-specific or fine-tuned variants, which may affect generalizability. Additionally, the use of LLMs to pick from multiple-choice options introduces potential selection bias, as highlighted by recent work Zheng et al. (2023). While we mitigate this by randomizing option order, it may not fully eliminate bias, impacting generation reliability.
Furthermore, our reliance on GPT-4o-mini as the primary evaluation judge could be questioned, as automated LLM-based scoring may introduce biases. Although we conducted a small-scale human evaluation to validate its reliability, finding substantial agreement with human annotators (Cohen’s kappa = 0.82), minor inconsistencies were observed in ambiguous cases (see Appendix B for details).
Appendix A Custom Prompts
We generate the answers for $\mathcal{N}$ setting using the following prompt. The prompt outputs the answer and confidence for the answer in a json format.
1LLM_RESPONSE_PROMPT = """
2You are an intelligent assistant who is given a question. Your role is to provide accurate, helpful, and well-reasoned responses based on your knowledge and capabilities.
3
4Along with the question, you need to provide a confidence score for your answer. The confidence score should be a number between 0 and 100, where:
5- 0-25 indicates low confidence
6- 26-75 indicates moderate confidence
7- 76-100 indicates high confidence
8
9Guidelines for providing answers:
101. Be direct and concise in your answer while ensuring completeness. Avoid unnecessary words or tangents.
112. If you are uncertain, provide a lower confidence score.
123. Base your confidence score on:
13 - The reliability and recency of available information
14 - Your knowledge of the specific domain
15
16Here are some examples:
17
18Example 1:
19Question: What is the capital of France?
20Answer: Paris
21Confidence score: 91
22(High confidence as this is a well-established fact)
23
24Example 2:
25Question: Which country has the best healthcare system?
26Answer: It depends on the criteria used. Some rankings favor Switzerland, while others favor Sweden or Singapore.
27Confidence score: 25
28(There is no definitive answer, and the confidence is low due to the lack of a clear consensus.)
29
30Example 3:
31Question: Which state is between Washington and California?
32Answer: Oregon
33Confidence score: 87
34(Maximum confidence as this is a clear geographic fact)
35
36Example 4:
37Question: What was Albert Einstein’s favorite food?
38Answer: There is no definitive record of his favorite food, but he reportedly liked pasta.
39Confidence score: 25
40(There are anecdotal mentions, but no verified records.)
41
42Example 5:
43Question: Is Irvine a city in California?
44Answer: Yes
45Confidence score: 81
46(High confidence as this is a verifiable fact)
47
48Example 6:
49Question: What is the most popular programming language for AI development?
50Answer: Python
51Confidence score: 66
52(Moderate-high confidence based on current trends, but this can change over time)
53
54Here is a new example. Simply reply with your answer and confidence score.
55
56Question: {{question}}
57
58Provide your response in the following JSON format:
59{
60 "answer": "Your answer here",
61 "confidence_score": number between 0-100
62}
63"""
We generate the answers for $\mathcal{D}$ setting using the following prompt. The prompt outputs the answer and confidence for the answer in a json format.
1LLM_RESPONSE_PROMPT_DISTRACTORS = """
2You are an intelligent assistant who is given a question and a list of options. Your role is to provide accurate, helpful, and well-reasoned answer based on your knowledge and capabilities and the options provided.
3
4Along with the answer, you need to provide a confidence score for your answer. The confidence score should be a number between 0 and 100, where:
5- 0-25 indicates low confidence
6- 26-75 indicates moderate confidence
7- 76-100 indicates high confidence
8
9Guidelines for providing answers:
101. Return the answer from the list of options provided only. It is guaranteed that the answer will be one of the options provided.
112. If you are uncertain, provide a lower confidence score.
123. Base your confidence score on:
13 - The reliability and recency of available information
14 - Your knowledge of the specific domain
15
16Here are some examples:
17
18Example 1:
19Question: What is the capital of France?
20Options:
21- Paris
22- London
23- Rome
24- Madrid
25Answer: Paris
26Confidence score: 91
27(High confidence as this is a well-established fact)
28
29Example 2:
30Question: Which country has the best healthcare system?
31Options:
32- Switzerland
33- Sweden
34- Singapore
35- United States
36Answer: It depends on the criteria used. Some rankings favor Switzerland, while others favor Sweden or Singapore.
37Confidence score: 25
38(There is no definitive answer, and the confidence is low due to the lack of a clear consensus.)
39
40Example 3:
41Question: Which state is between Washington and California?
42Options:
43- Oregon
44- Washington
45- California
46- Idaho
47Answer: Oregon
48Confidence score: 87
49(Maximum confidence as this is a clear geographic fact)
50
51Example 4:
52Question: What was Albert Einstein’s favorite food?
53Options:
54- Pizza
55- Pasta
56- Sushi
57- Tacos
58Answer: There is no definitive record of his favorite food, but he reportedly liked pasta.
59Confidence score: 25
60(There are anecdotal mentions, but no verified records.)
61
62Example 5:
63Question: Is Irvine a city in California?
64Options:
65- Yes
66- No
67Answer: Yes
68Confidence score: 81
69(High confidence as this is a verifiable fact)
70
71Example 6:
72Question: What is the most popular programming language for AI development?
73Options:
74- Python
75- Java
76- C++
77- JavaScript
78Answer: Python
79Confidence score: 66
80(Moderate-high confidence based on current trends, but this can change over time)
81
82Here is a new example. Simply reply with your answer and confidence score.
83
84Question: {{question}}
85Options: {{options}}
86
87Provide your response in the following JSON format:
88{
89 "answer": "Your answer here",
90 "confidence_score": number between 0-100
91}
92"""
To generate the 3 distractors for each question-answer pair, we use the following prompt with few shot examples.
1DISTRACTORS_GENERATION_PROMPT = """You are an expert synthetic data generator. Your task is to generate three plausible but incorrect answers to a given question.
2
3Guidelines for generating wrong answers:
41. Each answer should be factually incorrect but plausible within the context
52. Match the answer type (e.g. if asking for a date, provide wrong dates)
63. The wrong answers should be clearly distinct from the correct answer and from each other
74. Maintain a similar level of specificity as the original answer
85. The answers should be realistic and not obviously wrong
9
10Example 1:
11Question: What is the capital of France?
12Answer: Paris
13Wrong Answers:
14- Lyon
15- Marseille
16- Bordeaux
17Reason: All are major French cities, but incorrect as capital
18
19Example 2:
20Question: Who was the first president of the United States?
21Answer: George Washington
22Wrong Answers:
23- John Adams
24- Thomas Jefferson
25- Benjamin Franklin
26Reason: All are founding fathers but not the first president
27
28Example 3:
29Question: In what year did World War II end?
30Answer: 1945
31Wrong Answers:
32- 1943
33- 1944
34- 1946
35Reason: All are plausible years during or near WWII but not when it ended
36
37Example 4:
38Question: Who wrote Romeo and Juliet?
39Answer: William Shakespeare
40Wrong Answers:
41- Christopher Marlowe
42- Ben Jonson
43- John Webster
44Reason: All are prominent Elizabethan playwrights
45
46Example 5:
47Question: What is the largest planet in our solar system?
48Answer: Jupiter
49Wrong Answers:
50- Saturn
51- Neptune
52- Uranus
53Reason: All are gas giant planets, but smaller than Jupiter
54
55Please generate three wrong answers that follow these guidelines for the given question.
56The answers should be:
57- Factually incorrect but plausible
58- Match the same answer type (e.g. date, person, number)
59- Clearly distinct from the correct answer and each other
60- Similar in specificity/detail level
61- Realistic and not obviously wrong
62
63Return only three wrong answers as a list in JSON format with the following requirements:
64- Each wrong answer should be a string
65- The output should be a single JSON object with key "wrong_answers"
66- The value should be an array of exactly 3 wrong answers
67- No explanations or additional text should be included
68- The answers should maintain consistent formatting with the correct answer
69
70Example format:
71{{
72 "wrong_answers": ["opt1", "opt2", "opt3"]
73}}
74
75Question: {question}
76Correct Answer: {answer}
77Generate three wrong answers:
78"""
Table 3: Performance metrics of LLMs in the Normal ($\mathcal{N}$) and Distractor ($\mathcal{D}$) settings, including accuracy (correct), non-attempt (na), ECE, and the number of helped ($\mathcal{D}_{helped}$) and harmed ($\mathcal{D}_{harmed}$) instances.
LLMs $\mathcal{N}_{correct}$ $\mathcal{N}_{none}$ $\mathcal{N}_{ECE}$ $\mathcal{D}_{correct}$ $\mathcal{D}_{none}$ $\mathcal{D}_{ECE}$ $\mathcal{D}_{helped}$ $\mathcal{D}_{harmed}$
GPT-4o-mini [Uncaptioned image] 8.46% 6.80% 0.750 47.43% 0.02% 0.320 1644 (93.78%) 109
GPT-4-turbo [Uncaptioned image] 20.99% 7.14% 0.616 65.33% 0.02% 0.165 1821 (95.44%) 87
GPT-4o [Uncaptioned image] 36.75% 8.16% 0.437 73.48% 0% 0.037 1507 (91.22%) 145
LLaMA3.1-8b-instant [Uncaptioned image] 8.24% 19.58% 0.780 44.94% 0.21% 0.367 1294 (91.45%) 121
LLaMA 3-8B-8192 [Uncaptioned image] 9.27% 24.99% 0.790 45.56% 2.45% 0.355 1251 (90.46%) 132
Gemma2-9B-it [Uncaptioned image] 9.52% 34.49% 0.771 46.58% 1.87% 0.359 1060 (88.70%) 135
Appendix B Same LLM judge as Prediction LLM
We employ the same LLM as both the judge and the base model responsible for predicting answers to the questions. The performance metrics are presented in Table 3. Upon manually inspecting instances from smaller LLMs, we observe that the LLM judge occasionally misclassifies responses and refrains from assigning NOT_ATTEMPTED to certain data points. This can be seen by comparing $\mathcal{N}_{none}$ in Table1 and 3. To address this issue and ensure consistency, we use gpt-4o-mini as the LLM judge across all models. We also create reliability diagrams of pipeline where LLM-judge was different and the graphs are shown in Figure 4.
To validate the reliability of GPT-4o-mini as an LLM judge, we conducted a small-scale human evaluation. We sampled 100 responses and had three human annotators independently classify them as CORRECT, INCORRECT, or NOT_ATTEMPTED. The inter-annotator agreement, measured using Cohen’s kappa, was 0.82 for GPT-4o-mini, indicating substantial agreement. This comparison allowed us to measure the extent of biases introduced by automated evaluation and confirm that LLM judges generally aligned with human judgments, though minor inconsistencies were observed in ambiguous cases.
Figure 4: Reliability diagrams (RDs) showing calibration performance in $\mathcal{N}$ ($\bullet$) and $\mathcal{D}$ ($\bullet$) settings. The numbers on top of bars represent the number of correctly predicted instances. (y-axis: actual accuracy, x-axis: predicted confidence)
Appendix C LLMs details
In our study, we utilize six distinct LLMs representing a diverse range of capabilities and architectures. The GPT models (GPT-4o-mini, GPT-4-turbo, and GPT-4o) are state-of-the-art LLMs provided via OpenAI’s API. These models are known for their advanced reasoning, contextual understanding, and robust performance across a variety of tasks. The larger models, such as GPT-4o, excel in handling complex queries and demonstrate superior accuracy, while the smaller variant, GPT-4o-mini, serves as a lightweight alternative with reduced computational overhead. We use GroqCloud for the LLaMA3.1-8B-instant and LLaMA3-8B-8192 models. These models are smaller in scale but are effective in tasks requiring efficient knowledge retrieval and basic reasoning. Finally, Gemma2-9B-it, also from GroqCloud, offers a balance between size and capability, though its performance is limited by its relatively smaller model size and less extensive pretraining. Together, these models provide a comprehensive landscape for evaluating calibration, accuracy, and decision-making across diverse architectures and computational setups.